Priscilla Lane and Margaret Green studied the link
Answer #1

Answer:

A).

Order of genes = Rg—a---mg

Gene map:

Rg----------25.35mu--------a-----12.71mu-----mg

Coefficient of Coincidence (COC) = 0.29

Interference = 0.71

B). Triple heterozygote = + a mg / Rg + +

Explanation:

Hint: Always parental combinations are more in number than the any type of recombinants. Triple heterozygote = a + mg / + Rg +

1).

If single crossover occurs between a & rg

Normal combination: a + / + rg

After crossover: a rg / + +

a rg progeny= 1+36=37

+ + progeny = 1+16 = 17

Total this progeny = 54

Total progeny = 213

The recombination frequency between a & rg = (number of recombinants/Total progeny) 100

RF = (54/213)100 = 25.35%

2).

If single crossover occurs between rg & mg

Normal combination: + mg / Rg +

After crossover: + + / Rg mg

+ + progeny= 15+16=31

Rg mg progeny = 9+36= 45

Total this progeny = 76

The recombination frequency between rg & mg = (number of recombinants/Total progeny) 100

RF = (76/213)100 = 35.68%

3).

If single crossover occurs between a & mg

Normal combination: a mg / + +

After crossover: a + / + mg

a + progeny= 15+1=16

+ mg progeny = 1+9=10

Total this progeny = 26

The recombination frequency between a & mg = (number of recombinants/Total progeny) 100

RF = (26/213)100 = 12.71%

Recombination frequency (%) = Distance between the genes (mu)

Rg----------25.35mu--------a-----12.71mu-----mg

Expected double crossover frequency = (RF between Rg & a) * (RF between a & mg)

= 25.35% * 12.71% = 0.0322

The observed double crossover frequency = 1+1/213 = 0.0094

Coefficient of Coincidence (COC) = Observed double crossover frequency / Expected double crossover frequency

= 0.0094 / 0.0322

= 0.29

Interference = 1-COC

= 1-0.29 = 0.71